### v2

# Self-Translate-Train: Enhancing Cross-Lingual Transfer of Large Language Models via Inherent Capability

Abstract
Zero-shot cross-lingual transfer by fine-tuning multilingual pretrained models shows promise for low-resource languages, but often suffers from misalignment of internal representations between languages. We hypothesize that even when the model cannot generalize across languages effectively in fine-tuning, it still captures cross-lingual correspondence useful for cross-lingual transfer. We explore this hypothesis with Self-Translate-Train, a method that lets large language models (LLMs) to translate training data into the target language and fine-tunes the model on its own generated data. By demonstrating that Self-Translate-Train outperforms zero-shot transfer, we encourage further exploration of better methods to elicit cross-lingual capabilities of LLMs.
Self-Translate-Train: Enhancing Cross-Lingual Transfer
of Large Language Models via Inherent Capability
Ryokan Ri and Shun Kiyono and Sho Takase SB Intuitions {ryokan.ri,shun.kiyono,sho.takase}@sbintuitions.co.jp
1 Introduction
Cross-lingual transfer is a technique to solve tasks in a target language by leveraging data from other languages (Pikuliak et al., 2021). A popular approach is to fine-tune a multilingual pre-trained model (Conneau et al., 2020; Xue et al., 2021; Scao et al., 2022) on a source language data (e.g., English) and then apply it to the target language. Capable models can solve tasks in the target language without seeing its data during fine-tuning, which is known as zero-shot cross-lingual transfer (Artetxe and Schwenk, 2019; Chen et al., 2021).
However, zero-shot cross-lingual transfer is often suboptimal, with a performance gap between source and target languages. One potential factor for this gap is representation misalignment between languages (Artetxe et al., 2023): the model does not capture the correspondence between languages effectively. We hypothesize that in some cases the model does capture the correspondence, but just cannot utilize it when fine-tuned only on the source language data; we need to elicit its cross-lingual capability in a different way.
To explore this hypothesis, we propose and investigate an approach called Self-Translate-Train, which leverages the powerful text generation capabilities of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023). Instead of fine-tuning an LLM on source language data and hoping for generalization, Self-Translate-Train first let the model translate the training data into the target language, and fine-tunes the model on the data that the model itself generates. This process does not only rely on cross-lingual generalization that happens during fine-tuning, but also on the cross-lingual correspondence that the model captures and can be elicited through translation.
This approach is inspired by the Translate-Train (Conneau et al., 2018; Hu et al., 2020; Artetxe et al., 2023) approach, a method to translate training data to the target language, often with an external machine translation system. Notably, our findings reveal that the Translate-Train approach remains effective even without external resources.
We validate the effectiveness of Self-Translate-Train with three tasks (math, QA, NLI) across four target languages (de, ru, th, zh). The results show that better cross-lingual performance can be achieved by correctly eliciting the model’s translation capability, which encourages further exploration of a better method for cross-lingual transfer.
2 Related Work
2.1 Cross-lingual Transfer Learning
There are two main approaches to transfer task knowledge across languages: model transfer and data transfer (Pikuliak et al., 2021).
Model transfer leverages multilingual pre-trained models. As the model is expected to capture the commonality between languages, when fine-tuned on source language data, it can generalize to solve the task in other languages (Pires et al., 2019; Mulcaire et al., 2019; Conneau et al., 2020).
Data transfer translates the source language data into the target language. In the Translate-Test approach, models are assumed to be specialized in the source language and the task inputs are translated into this language at test time (Conneau et al., 2018; Etxaniz et al., 2024). The Translate-Train approach translates the training data, and the trained model makes predictions in the target language (Conneau et al., 2018). Data transfer is often more effective than model transfer (Hu et al., 2020; Artetxe et al., 2023; Ebing and Glavaš, 2024).
Our approach, Self-Translate-Train, directly translates the training data as in data transfer, but also leverages the cross-lingual capabilities of the model itself as in model transfer.
2.2 Self-Improvement of LLMs
LLMs have remarkable text generation capabilities, which has been leveraged to generate training data for various purposes (Li et al., 2023b; Lee et al., 2024). The generated data can be used to further specialize the LLM itself for downstream applications, without requiring an extensive collection of additional data. This process can be viewed as a form of self-improvement (Bai et al., 2022; Huang et al., 2023; Sun et al., 2023; Li et al., 2023a).
Self-Translate-Train is also a self-improvement approach to specialize the LLM to a target language by translating the source language data.
3 Self-Translate-Train
Our framework focuses on fine-tuning LLMs on a small amount of data for a specific task. Let the training corpus in a source language, say English, be , where is the input and is the output. In a typical cross-lingual transfer setting, the model is fine-tuned only on and expected to generalize to target languages.
Translated Synthetic Data
Given the LLM’s translation capability, we can let it translate the training corpus into a synthetic corpus in the target language . The synthetic data can be added to to achieve a better generalization to the target language.
The translation can be performed in various ways depending on the model’s capabilities or available resources. In this paper, we experiment with the few-shot prompting technique (Section 4.4).
Code-switched Synthetic Data
The generated data has an interesting aspect: each synthetic instance has a corresponding instance in the original dataset with the same semantics. We can exploit this to further synthesize data by generating code-switched instances where the input and output are in different languages.
We pair the original and translated instances to construct . When the task output is natural language, we manually translate the prompt “Please answer in {{ tgt }}.” into the target language, and add it to the input .
4 Experimental Setups
To verify the effectiveness of Self-Translate-Train, we conduct extensive experiments on multiple tasks and languages.
4.1 Task and Datasets
We present a list of datasets for experiments in Table 1. For each task, an English dataset is used for training and a multilingual dataset for evaluation. To make the computational cost feasible, we use a 10,000-sample subset of the training data for SQuAD and MultiNLI.
| Task | Training | Evaluation |
| QA | SQuAD (Rajpurkar et al., 2016) | XQuAD (Artetxe et al., 2020) |
| Classification | MultiNLI (Williams et al., 2018) | XNLI (Conneau et al., 2018) |
| Math | GSM8k (Cobbe et al., 2021) | MGSM (Shi et al., 2023) |
4.2 Languages
We conducted evaluation on four languages: German (de), Russian (ru), Thai (th), and Chinese (zh). German is a Germanic language, which is phylogenetically close to English and expected to show better cross-lingual transfer, while Russian, Thai, and Chinese are from different language families. In particular, Thai is a low-resource language with a different script from English, which is expected to be more challenging for cross-lingual transfer.
4.3 Language Models
4.4 Synthetic Data Generation
Recent LLMs are known to exhibit a translation capability without much task-specific data Briakou et al. (2023). In our experiments, we elicit the translation capability of the LLMs via few-shot in-context learning (Brown et al., 2020).
To construct few-shot translation samples, we sample eight pairs from the train or validation splits of the multilingual datasets, where instances across languages form parallel data. The translation was performed for each field individually, e.g., for GSM8k, we translated the question and answer separately. The prompt template simply alternates the source and target text prepended with the language tag (Section A.2).
An important step to ensure the quality of the synthetic data is to filter out the low-quality data (the details in Section A.3). To remove under- or over-translation (Tu et al., 2016), we filter out texts with an extreme source-target length ratio. Also, to address the repetition problem (Holtzman et al., 2020), we set the max number of tokens for generation and filter out the translation that does not end with the EOS token. With the translations from Llama2-7B, this process removes around 10% of the data for most languages and around 50% for Thai due to the model’s limited generation quality.
To provide the sense of the translation quality, we report the BLEU score (Papineni et al., 2002) measured by the parallel data constructed from questions in the MGSM test set in Table 2. Overall, the translation quality is sufficiently high except for Thai. As we will see in Section 5.1, this poses a challenge for cross-lingual transfer to Thai.
| Model | de | ru | th | zh |
| Llama2-7B | 37.1 | 27.2 | 1.9 | 29.4 |
4.5 Fine-tuning
All the tasks are cast as text generation tasks, where the LLM is given the inputs as a prompt and generate the answer. Fine-tuning is conducted with causal language modeling loss, computed only for output tokens. We use LoRA (Hu et al., 2022), a parameter-efficient tuning technique, to reduce computational cost.
We use AdamW (Loshchilov and Hutter, 2019) and the cosine learning rate schedule for optimization, training with a batch size of 64 for 1,000 steps. For each setting, we conduct six runs with two learning rates (5e-5 and 3e-4) and different random seeds, reporting summarization statistics of the top four runs based on validation accuracy to remove runs with optimization failure. See Section A.5 for other hyperparameters.
5 Results
5.1 Main Results
| MGSM | XQuAD | XNLI | ||||||||||||||||||||||||||||||||||
| de | ru | th | zh | de | ru | th | zh | de | ru | th | zh | |||||||||||||||||||||||||
| FewShot |
As a baseline, we perform few-shot learning with the target language data with the pre-trained model (FewShot). We also fine-tune the LLM with the source language dataset . To ensure equal data resources, we augment with the eight target language samples used for few-shot translation (Section 4.4). We then compare these baselines with the models fine-tuned on the data generated from Self-Translate-Train ( and ) in Table 3111For experiments with other approaches, including translate-test and translation with external systems, see Appendix C..
Results show that fine-tuning on English () is more effective than few-shot learning in the target language (FewShot). This confirms that cross-lingual transfer from English is non-trivial and effective in this setting.
Notably, Self-Translate-Train is indeed an effective method; almost consistently outperforms the baseline . The only exception is Thai (th), where there is no significant improvement. This is likely due to the low translation quality of the model in Thai (Table 2).
However, the effectiveness of code-switching dataset is limited. When we add to , there is no significant improvement from adding only ( in Welch’s t-test). This indicates that the code-switching data does not provide additional information for the model to generalize in the task.
5.2 Does the model size matter?
The size of the LLM can influence both its ability to generalize across languages and the quality of its translations, which in turn may impact the effectiveness of Self-Translate-Train. We compare the performance of Llama2 with different sizes, i.e., 7B, 13B, and 65B, on the math task in de, th, and zh. For each model size, we generate with the corresponding model and then fine-tune itself.
Figure 2 shows that the larger model size generally tends to perform better, and the improvement from Self-Translate-Train remains consistent across different model sizes. In Thai (th), we did not observe a significant improvement in the 7B model, but do in the larger models (13B and 65B), likely due to their better translation quality. The 7B model has a low Thai translation BLEU score of 1.9 (Table 2), while the 13B and 65B models have BLEU scores of 5.1 and 12.0, respectively.
The improvement in Thai (th) with the 70B model is the most significant (+19.8 average points). This implies that Self-Translate-Train is particularly effective when the model struggles with generalizing across the source and target languages but can still generate their reasonable translations.
6 Discussion and Conclusion
We explored Self-Translate-Train, a method to improve cross-lingual transfer performance by generating synthetic training data in the target language. We validated its effectiveness on various tasks and languages, demonstrating substantial performance gains over baselines with the same data budget.
This result aligns with Etxaniz et al. (2024), who showed that the performance of few-shot learning can be improved by self-translating the task inputs. Our work demonstrates that self-translation by an LLM is also effective when applied to task-specific training data for fine-tuning.
These results suggest that we can improve the cross-lingual transfer performance of an LLM without any additional data, but by investing more computation to generate synthetic data. This finding is consistent with the recent trend of self-improvement methods for LLMs (e.g., Huang et al., 2023).
Future work could explore synthetic data generation for pre-training to improve cross-lingual capabilities of LLMs without intensive data collection. We hope that this work encourages further exploration of how to better elicit models’ cross-lingual capabilities.
7 Limitations
Our experiments are conducted on a modern type of LLM, an autoregressive Transformer decoder, and centered around the Llama2 model families (Touvron et al., 2023). Although we further validate our method in Appendix B, the effective of our proposed method is uncertain when applied to other types of LLMs developed in the future.
Our method is based on the assumption that the model can generate reasonable translations in the target language. However, for the low-resource target language or the datasets with out-of-domain text for the model, the translation quality can be not satisfactory for the Translate-Train approach (Ebrahimi and von der Wense, 2024).
Finally, when the task requires generating long and natural text, the quality of the generated translation matters more. If the translation quality is low, the model outputs may degrade due to translation errors or unnaturalness. The application of our method on more challenging tasks requires further technical considerations.
Appendix A The Details of Experimental Setups
A.1 Tasks and Datasets
We provide the details of the datasets introduced in Section 4.1.
A.1.1 Question Answering (QA)
SQuAD (Rajpurkar et al., 2016) is an English QA dataset created from Wikipedia articles as training data. Given a question and a passage, the task is to extract the answer from the passage. Evaluation is conducted with XQuAD (Artetxe et al., 2020), which consists of translation of SQuAD into multiple languages.
Context: Architecturally, the school has a Catholic character. Atop the Main Building’s gold dome is a golden statue of the Virgin Mary. Immediately in front of the Main Building and facing it, is a copper statue of Christ with arms upraised with the legend "Venite Ad Me Omnes". Next to the Main Building is the Basilica of the Sacred Heart. Immediately behind the basilica is the Grotto, a Marian place of prayer and reflection. It is a replica of the grotto at Lourdes, France where the Virgin Mary reputedly appeared to Saint Bernadette Soubirous in 1858. At the end of the main drive (and in a direct line that connects through 3 statues and the Gold Dome), is a simple, modern stone statue of Mary. Question: To whom did the Virgin Mary allegedly appear in 1858 in Lourdes France? |
Saint Bernadette Soubirous |
A.1.2 Text-Pair Classification
We also evaluate our method on cross-lingual text-pair classification tasks. The MultiNLI dataset (Williams et al., 2018) involves determining the logical relationship between a premise sentence and a hypothesis sentence. XNLI (Conneau et al., 2018) is a multilingual NLI dataset for evaluation.
Premise: Conceptually cream skimming has two basic dimensions - product and geography. Hypothesis: Product and geography are what make cream skimming work. What is their logical relation? Entailment, Neutral or Contradition. |
Neutral |
A.1.3 Mathematical Reasoning
GSM8k (Cobbe et al., 2021) is an English dataset of 8.5K high-quality grade school math problems. Each problem is annotated with a solution that shows the mathematical steps required to reach the final answer. As the evaluation dataset, we use MGSM (Shi et al., 2023), a multilingual version of the GSM8k dataset.
The LLM is trained to generate the step-by-step solution to math problems. The answer is extracted from the LLM output as the final digits, and the accuracy is calculated based on the exact match of the extracted answer and the ground truth.
Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May? |
Natalia sold 48/2 = clips in May. Natalia sold 48+24 = 72 clips altogether in April and May. #### 72 |
A.2 Prompt Format for LLM Translation
To translate training data using a LLM (Section 4.4), we employed the following prompt template for each task. The template simply consists of the source text and target text prepended with the language tag. The text is surrounded by backticks and the LLM starts generating the target text the open backtick until the close backtick is found.
{% for sample in few_shot_samples %}
en: ‘{{ sample.data_field }}‘
{{ target_language }}: ‘{{ sample.data_field }}‘
{% endfor %}
en: ‘{{ data_field }}‘
{{ target_language }}: ‘
The SQuAD dataset annotates answer spans in the context passages. We translate the annotations using the mark-then-translate approach (Chen et al., 2023). We mark the answer span in the context passage with the tokens “<answer>” and “</answer>”, translate the marked text, and then extract the translated answer span from the translated context. Note that in this case, the few-shot samples are also marked with the answer span.
The context passages in the SQuAD dataset are relatively long, and it is challenging for the LLM with a limited context window to fit the entire few-shot samples and the source text. To address this issue, we split the context into sentences using the spaCy library222https://spacy.io/ and translate them separately, i.e., the few-shot samples and the source text are sentences.
A.3 Data Filtering for Synthetic Data
We remove pairs where the target length is less than one-third or more than three times the source length. The text length is heuristically determined to account for character length differences between languages. For example, phonogram-based text (e.g., English) has much more characters than ideograph-based text (e.g., Chinese). We set normalization factors where English, German, Thai, and Russian characters count as 1, and Chinese characters as 3.
We also filter out incomplete translations which are typically produced by repetitive generation. We set the maximum number of tokens for generation (Table 5) and remove the outputs not ending with the token indicating the end of the translation, in our case, the backtick character used in the prompt format.
A.4 Assessing the Translation Quality
To evaluate the translation quality, we use the BLEU score (Papineni et al., 2002) measured by the parallel data constructed from questions in the MGSM test set. The translation is performed in few-shot in-context learning with 8 translation samples constructed from the train set of the MGSM dataset. The BLEU score is calculated using the SacreBLEU library (Post, 2018)333https://github.com/mjpost/sacrebleu. As the tokenizer option, we use “13a” for de and ru, “flores101” for th, and “zh” for zh.
Table 6 shows the BLEU scores from the LLMs evaluated in this paper. The result of Qwen1.5-1.8B is discussed in Appendix B, and gpt-3.5-turbo-0125 in Section C.4.
| Model | de | ru | th | zh |
| Llama2-7B | 37.1 | 27.2 | 1.9 | 29.4 |
| Llama2-13B | 41.3 | 33.4 | 5.1 | 34.3 |
| Llama2-70B | 45.6 | 41.9 | 12.0 | 42.4 |
| Qwen1.5-1.8B | 21.9 | 11.3 | 1.5 | 41.3 |
| gpt-3.5-turbo-0125 | 48.0 | 44.6 | 23.1 | 47.2 |
A.5 Hyper-parameters for Fine-tuning
We provide the hyper-parameters used for fine-tuning the LLMs in Table 7.
| Hyper-parameter | Value | Hyper-parameter | Value |
| Batch size | 64 | Adam | 1e-8 |
| Number of steps | 1,000 | Adam | 0.9 |
| Learning rate | [5e-5, 3e-4] | Adam | 0.999 |
| LR Scheduler | Cosine | Weight decay | 0.1 |
| Warmup ratio | 0.05 |
Appendix B Results from Qwen1.5-1.8B
To increase the robustness of the results, we also conducted experiments with Qwen1.5-1.8B 444https://qwenlm.github.io/blog/qwen1.5/. While the model is mainly trained on Chinese and English data, it is also constructed with the multilingual use cases in mind.
| MGSM | ||||||||||||
| de | ru | th | zh | |||||||||
We observe that the results are consistent with the main experiments: Self-Translate-Train is effective when the zero-shot cross-lingual transfer performance is suboptimal and the model can generate reasonable translations. The performance is improved by adding the target language data . However, when the translation quality is poor as in Thai (1.5 BLEU score in Table 6), the improvement is not observed. Additionally, Qwen1.5-1.8B seems to have good cross-lingual capability between English and Chinese, as indicated by the high BLEU score (41.3 in Table 6). With this, tuning on the source language data alone is sufficient to achieve high performance.
Appendix C Additional Exepriments
In this section, we discuss experiments that are outside the scope of the main topic of this paper but are somewhat relevant and may be of interest to readers.
C.1 Does Self-Translate-Train improve the performance in the source language?
The performance somtimes improves, given the task is challenging and the translation quality is sufficiently high.
Table 9 shows the results on the English test set with Llama2-7B. The performance improves in the MGSM dataset when adding the synthetic data from de, ru, and zh. The Thai language does not show the improvement possibly due to the low translation quality.
However, the improvement is not observed in the XQuAD and XNLI datasets. This might be because the task performance is already high with the source language data alone, and the synthetic data does not provide additional information to improve the performance.
| MGSM | XQuAD | XNLI | ||||||||||||||||||||||||||||||||||
| de | ru | th | zh | de | ru | th | zh | de | ru | th | zh | |||||||||||||||||||||||||
C.2 Does the synthetic data alone improve the performance in the target language?
Yes, but adding the source language data is more effective. Table 10 shows the results with Llama2-7B on the multilingual evaluation datasets. Tuning on the synthetic data alone () improves the performance in the target language, but the improvement is not as significant as adding the synthetic data to the source language data (). In practice, we recommend using the synthetic data in combination with the original data to achieve the best performance.
| MGSM | XQuAD | XNLI | ||||||||||||||||||||||||||||||||||
| de | ru | th | zh | de | ru | th | zh | de | ru | th | zh | |||||||||||||||||||||||||
C.3 What about the Translate-Test approach?
The translate-test approach, where non-English inputs are translated into English and fed into an English model (i.e., the model trained with ), can be an alternative to the trainslate-train approach. We perform experiments with the translate-test approach and provide additional context on the effectiveness of Self-Translate-Train. The tasks to use are the math and NLI tasks, whose outputs are language-independent 555We cannot simply apply the translate-test approach to QA, as the output language would be different from the original language. We could remedy this by further translating the output back to the original language (Asai et al., 2018), but this is out of the scope of our exploration here..
We first experiment with the Self-Translate-Test approach in few-shot learning settings (Etxaniz et al., 2024). This approach translates the test-time inputs to high-resource language (e.g., English) by few-shot translation and performs inference on the translated inputs.
We compare the Self-Translate-Test approach with a baseline without translation. The FewShot baseline uses 8-shot samples in the target language from the train set for inference on the test set. The Self-Translate-Test (few-shot) approach translates the test set to English and performs inference using English 8-shot samples. The results are shown in Table 11.
| MGSM | XNLI | |||||||||||||||||||||||
| de | ru | th | zh | de | ru | th | zh | |||||||||||||||||
| FewShot |
| Self-Translate-Test (few-shot) |
We can see that the Self-Translate-Test approach is effective both in the math and NLI tasks, which conforms to the results in Etxaniz et al. (2024).
Next, we validate the effectiveness in fine-tuning settings. Specifically, we feed the model fine-tuned with English task-specific data with test inputs translated into English by few-shot translation of the pre-trained model. The results are shown in Table 12.
| MGSM | XNLI | |||||||||||||||||||||||
| de | ru | th | zh | de | ru | th | zh | |||||||||||||||||
| + Self-Translate-Test |
We observe that the Self-Translate-Test approach remains effective when combined with a fine-tuned model. This strengthen the argument that the self-translate approach is an effective way to improve the cross-lingual performance of the model under the same data budget. However, this approach is not always practical, as it requires both the pre-trained model for few-shot translation and the fine-tuned model to be available during deployment.
C.4 How effective is generated by an external system?
When the synthetic training data is generated by an external system, such approach can be seen as the Translate-Train approach (Section 2.1) or sequence distillation from a teacher model (Kim and Rush, 2016). As an upper-bound experiment using the math task, we fine-tune Llama2-7B on the synthetic data generated by gpt-3.5-turbo-0125 from the OpenAI API666https://openai.com/index/openai-api/, which produces high-quality translations across the languages explored in this paper (Table 6).
| MGSM | ||||||||||||
| de | ru | th | zh | |||||||||
| from Llama2-7B |
| from Llama2-7B |
| from gpt-3.5-turbo-0125 |
| from gpt-3.5-turbo-0125 |
The synthetic data generated by gpt-3.5-turbo-0125 improves performance across languages over the baseline and the synthetic data generated by Llama2-7B. This confirms that the quality of the synthetic data is crucial for the performance improvement.
Note that the outputs from external models are often restricted in its usage777For example, Term of use of OpenAI API (January 31, 2024) restricts the usage of the outputs for training a model that competes with the API (https://openai.com/policies/terms-of-use/). Meta Llama 3 License (April 18, 2024) prohibits using the outputs to improve any other large language model (https://llama.meta.com/llama3/license/)., while the method explored in this paper can be used with the model at hand without other resources. Additionally, our interest in this paper is rather to explore the cross-lingual potential of the model itself and how to better utilize it.